# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ 2.1 Postulates of Information
If $f(X)\underline{\triangleright}Y$, we also denote this as $f^{-1}(y)\underline{\triangleleft}x$ or $x\underline{\triangleright}f^{-1}(y)$, to maintain consistency in the directionality of enabling.
Based on the concept of enabling mapping, Xu2024 proposed four postulates of information:
Postulate 1 (Binary Attribute) Every piece of information comprises two components: the ontology, which represents the intrinsic essence of the information, and the carrier, which embodies the objective form of the information.
Postulate 2 (Existence Duration) The content and form of information have their own respective temporal existences.
Postulate 3 (State Representation) The states of the ontology and the carrier of information have their own respective representations.
Postulate 4 (Enabling Mapping) The state of the ontology can be mapped onto the state of the carrier via an enabling mapping.
These four postulates are concise, align with common conceptions of information, and are compatible with the basic concepts of major branches of information science. Their scientific significance is reflected in:
  • Unification: Providing a common foundation for long-dispersed theories in information science;
  • Axiomatization: Allowing the definition and reasoning of information to be axiomatized and formalized, akin to mathematics and physics;
  • Universality: Facilitating application to cutting-edge fields such as the Internet, big data, artificial intelligence, cognitive science, and cyber-physical systems.
The concept of "enabling mapping" is novel. It requires not only that a surjective mapping from the ontological state to the carrier state be established mathematically, but also that the existence of the ontological state renders the carrier state an objective reality, highlighting the dual character of information as both a mathematical mapping and a physical realizationShannon1948 ; Landauer1996 ; Wiener2019 .
2.2 Formal Representation of States
Definition 2 (Higher-Order Formal Logical System): Suppose a formal system $\mathcal{L}$ is described using the following symbol set:
Variables $x_{1},x_{2},\ldots$;
Individual constants $a_{1},a_{2},\ldots$;
Functions $f_{1}^{1},f_{2}^{1},\ldots,f_{1}^{2},f_{2}^{2},\ldots,f_{1}^{3},f_{2}^{3},\ldots$;
Parentheses $(,)$;
Predicates $A_{1}^{1},A_{2}^{1},\ldots,A_{1}^{2},A_{2}^{2},\ldots,A_{1}^{3},A_{2}^{3},\ldots$;
Logical connectives "not" $\sim$, "implication" $\rightarrow$;
Quantifier "for all" $\forall$;
A term in $\mathcal{L}$ is defined as follows:
(1) Variables and individual constants are terms;
(2) If $f_{i}^{n}(n>0,i>0)$ is a function in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $f_{i}^{n}$ itself, then $f_{i}^{n}(t_{1},\ldots,t_{n})$ is also a term in $\mathcal{L}$;
(3) If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$, and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ that do not contain $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is a term in $\mathcal{L}$.
(4) The set of all terms in $\mathcal{L}$ is generated by (1), (2), and (3).
Based on this, we can define atomic formulas in $\mathcal{L}$:
If $A_{i}^{n}(n>0,i>0)$ is a predicate in $\mathcal{L}$ and $t_{1},\ldots,t_{n}$ are terms in $\mathcal{L}$ not containing $A_{i}^{n}$ itself, then $A_{i}^{n}(t_{1},\ldots,t_{n})$ is an atomic formula in $\mathcal{L}$.
We then define well-formed formulas (WFFs) in $\mathcal{L}$:
(5) Every atomic formula in $\mathcal{L}$ is a WFF;
(6) If $\mathcal{A}$ and $\mathcal{B}$ are WFFs in $\mathcal{L}$, then $\sim\mathcal{A}$, $\mathcal{A}\rightarrow\mathcal{B}$, $(\forall x_{i})\mathcal{A}$, $(\forall f_{i}^{n})\mathcal{A}$, and $(\forall A_{i}^{n})\mathcal{A}$ are all WFFs in $\mathcal{L}$, where $x_{i}$, $f_{i}^{n}$, $A_{i}^{n}$ are variables, functions, and predicates in $\mathcal{L}$, respectively, and no predicate or function symbol is applied, either directly or indirectly, to itself.
(7) The set of all WFFs in $\mathcal{L}$ is generated by (5) and (6).
The higher-order formal logical system $\mathcal{L}$ defined here, based on the flexibility of predicate logic, can describe various relationships and constraints among objects in a unified way. Through higher-order quantification support, the nested definition of complex data and behaviors, and it explicitly prohibits anypredicate or function symbol from applying to itself—directly or indirectly—to avoid self-referential paradoxes.
For simplicity, $\mathcal{L}$ does not define connectives such as "and" $\land$, "or" $\lor$, "if and only if" $\leftrightarrow$, or the quantifier "exists" $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of symbols and can make many formulas more conciseHamilton1988 . Therefore, they are introduced as defined symbols in subsequent discussions.
Definition 3 (Interpretation of a Formal System) An interpretation $E$ of a formal system $\mathcal{L}$ over a domain $D_{E}$ assigns specific values to constants, elements of the domain $D_{E}$ to variables, and turns function and predicate symbols into functions and predicates over that domain $D_{E}$, thereby yielding a set of WFFs each with a truth value.
Definition 3 assigns concrete values and meaning to a set of formulas in a formal system $\mathcal{L}$ through interpretation, thereby providing a fundamental approach for formally describing the definitional, attributive, structural, relational, and evolutionary state properties of objects as both the subjective and objective worlds.
Postulate 5 (Expression of Object State):The state set of an object collection over a specified time set is constituted by its attributes—such as properties, structures, values, and relations—as well as the evolution rules and logical combinations among these attributes, and satisfies the following conditions:
Theorem 1 (Formal Expression of States) If the state set $S(X,T)$ of objects $X$ over times $T$ satisfies Postulate 5, then it can be represented as an interpretation of a formal system $\mathcal{L}$ over the domain $X\times T$, and it is logically coherent.
Proof: Since the state set $S(X,T)$ of the object set $X$ over the time set $T$ satisfies Postulate 5, we construct the corresponding formal system $\mathcal{L}$ as follows:
By (3), the functions and predicates in $\mathcal{L}$ represent the properties, forms, values, and relations of $X$ over $T$ as captured in $S(X,T)$. These properties can be expressed through functions and predicates in the formal system $\mathcal{L}$ that conform to Definition 1.
By (4), consider two ternary predicates $A_{1}^{3}$ and $A_{2}^{3}$.And then let $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ and $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$ be two properties of $S(X,T)$ corresponding to objects $x_{1},x_{2}\in X$, times $t_{1},t_{2}\in T(t_{1}<t_{2})$, and binary functions $f_{1}(x_{1},t_{1}),f_{2}(x_{1},t_{1})$. Suppose all $A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))$ satisfy the rule of evolving from $t_{1}$ to $t_{2}$ into some $A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2}))$. The formula $(\forall A_{1}^{3}\forall f_{1}^{2}\forall x_{1}A_{1}^{3}(x_{1},t_{1},f_{1}^{2}(x_{1},t_{1}))\rightarrow\exists A_{2}^{3}\exists f_{2}^{2}\exists x_{2}A_{2}^{3}(x_{2},t_{2},f_{2}^{2}(x_{2},t_{2})))$ is also a well-formed formula in $\mathcal{L}$, with respect to the interpretation of $x_{1},x_{2},t_{1},t_{2}$, and is able to express the corresponding rules in $S(X,T)$. For other rules in $S(X,T)$, the related elements in the domain $X\times T$ can similarly be interpreted and expressed via corresponding WFFs in $\mathcal{L}$, with no implicit logical self-reference.
Therefore, $S(X,T)$ constitutes an interpretation $E$ of $\mathcal{L}$ over the domain $X\times T$. By (5), both itself and its logical closure are logically consistent. The theorem is thus proved.
Note that Theorem 1 stipulates that variables in $\mathcal{L}$ are elements of $X\times T$; thus, in formulas, object and time elements appear in pairs $(x,t)$. In practice, some attributes may be time-invariant; then simplified expressions may be used to describe such attributes without time.
Theorem 1 indicates that, by employing the higher-order formal system $\mathcal{L}$, it is possible to describe the states of virtually all things in the world. Both theoretical and empirical analyses qiu2025 further demonstrate that $\mathcal{L}$ is broadly applicable to the expression of state in disciplines such as mathematics, economics, ecology, computer and information systems, and natural language. This provides a unified state modeling theory for interdisciplinary research, which is of significant importance to fields such as artificial intelligence, knowledge engineering, and complex systems.
2.3 Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata.
Theorem 2 (State Representation of Finite Automata): For any finite automaton $M$, one can describe its input, output, and state-transition behavior over a related time set $T$ by the state set $S(M,T)$.
Proof: Let $M=(Q,R,U,\delta,\lambda)$ be a finite automaton, and $T=\{t_{i}|i=1,\ldots,n\}$ the (strictly increasing) time points at which $M$ performs input, output, or state-transition actions. By finite automaton theory mathsea2002 , $Q=\{q_{t_{i}}|t_{i}=1,\ldots,n-1\}$, $R=\{r_{t_{i}}|t_{i}=2,\ldots,n\}$, and $U=\{u_{t_{i}}|t_{i}=1,\ldots,n\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
$\delta:U\times Q\to U$ is the next-state function;
$\lambda:U\times Q\to R$ is the output function.
For each state $u_{t_{i}}\in U$, define the state predicate $State^{3}$ by
$\varphi_{U}(t_{i})=State^{3}(M,t_{i},u_{t_{i}})$ (2.3)
to assert that at time $t_{i}$, $M$ is in state $u_{t_{i}}$$i=1,\ldots,n$.
For each input $q_{t_{i}}\in Q$, define the input predicate $Input^{3}$ by
$\varphi_{I}(t_{i})=Input^{3}(M,t_{i},q_{t_{i}})$ (2.4)
to assert that at time $t_{i}$, $M$ receives input $q_{t_{i}}$$i=1,\ldots,n-1$.
For each output $r_{t_{i}}\in R$, define the output predicate $Output^{3}$ by
$\varphi_{O}(t_{i})=Output^{3}(M,t_{i},r_{t_{i}})$ (2.5)
to assert that at time $t_{i}$, $M$ produces output $r_{t_{i}}$$i=2,\ldots,n$.
The transition function $\delta$ is captured by the well-formed formula
$\varphi_{\delta}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\delta(u_{t_{i}},q_{t_{i}})\land Eq^{2}(u_{t_{i+1}},\delta(u_{t_{i}},q_{t_{i}}))\land\varphi_{U}(t_{i+1})$ (2.6)
@@ 2.3 Finite Automaton States
Similarly, the output function $\lambda$ is given by
$\varphi_{\lambda}(t_{i})=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land Eq^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land$ (2.7)
$IsElementof^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$

### v10

@@ 3.1 Definition of Information
Corollary 1 If $f$ is an enabled mapping from $X$ to $Y$, then $|X|\geq|Y|$, where $|X|$ and $|Y|$ denote the cardinalities of the sets $X$ and $Y$, respectively.
If $f(X)\trianglerighteq Y$, we also denote this as $f^{-1}(y)\trianglelefteq x$ or $x\trianglerighteq f^{-1}(y)$, to maintain consistency in the directionality of enabling. Fig.1(a) intuitively illustrates the dual nature of an enabling mapping, encompassing both its mathematical mapping and physical enabling attributes.
Figure 1: Schematic Diagram of Enabling Mapping, Noisy Information, and Information Transmission Chain. (a) The scene of penguin activity $X$ is mapped by an enabling mapping $f$ to an image $Y$ on a laptop; between $X$ and $Y$ there exists both a mathematical mapping relationship and a causal relationship. Moreover, mapping $X$ to $Y$ requires corresponding physical conditions and the expenditure of a certain amount of energy. (b) The left panel depicts the ontological state of noiseless information, while the right panel shows a noisy information state: the set $S_{o}^{-}$ is removed from the noiseless ontological state, and the previously disjoint set $S_{o}^{+}$ is superimposed, resulting in a noisy ontological state. (c) In the penguin activity information transmission chain, the carrier state of any subsequent piece of information can be regarded as the result of an enabling mapping from any prior ontological information state, and can also be reasonably interpreted by means of this ontological state and the corresponding enabling mapping.
Based on this, a more rigorous theoretical foundation for the definition of information can be established:
Definition 2 (Mathematical Definition of Information) Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\mathbb{\subseteq O}\cup\mathbb{S}$, time of occurrence $T_{h}\mathbb{\subseteq T}$, ontological state set $S_{o}\left(o,T_{h}\right)$, carrier $c\mathbb{\subseteq O}$, reflection time $T_{m}\mathbb{\subseteq T}$, and carrier reflection set $S_{c}\left(c,T_{m}\right)$ are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}\left(o,T_{h}\right)$ to $S_{c}\left(c,T_{m}\right)$, succinctly denoted as $I=\left\langle o,T_{h},S_{c},c,T_{m},S_{o}\right\rangle$. The set of all information $I$ is called the information space, denoted by $\mathfrak{T}$, which, together with matter and energy, forms the three fundamental constituents of the objective world (Xu et al., 2025).
It should be noted that both $S_{o}\left(o,T_{h}\right)$ and $S_{c}\left(c,T_{m}\right)$ can each be represented as the interpretation of a set of WFFs within a formal system. However, the respective formal systems need not be identical, which constitutes one of the essential characteristics of information—its capacity to establish connections among all entities in the world, regardless of differences in their underlying formal systems.
Reference (Shannon, 1948) is regarded as a foundational work in information theory. The paper proposed that ” $H=-\sum{p_{i}\log p_{i}}$ plays a central role as the measure of information in information theory.” The discussion in this paper includes both the source and channel of information. The source $Is$ is the origin of the information, and it generates events $e_{1}$, $e_{2},$…, $e_{n}$ with probabilities $p_{1}$, $p_{2},$…, $p_{n}$, respectively. The channel $Ic$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $Is$, the channel $Ic$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$ (where $1\leq i\leq n,\ \tau>0$). We define the ternary predicate ${IsEvent}^{3}(Is,T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $Is$ at time $T_{h}$, and the ternary predicate ${IcCode}^{3}(Ic,T_{m},c_{i})$ to represent that the channel $Ic$ received and transmitted code $c_{i}$ in the time interval $T_{m}$. Then,
$S_{Is}(Is,T_{h})={IsEvent}^{3}(Is,T_{h},e_{i})$ (3.3)
and
$S_{Ic}(Ic,(T_{h},T_{h}+\tau))={IcCode}^{3}(Ic,(T_{h},T_{h}+\tau),c_{i})$ (3.4)
are the states of the source $Is$ at time $T_{h}$ and the state of the channel $Ic$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $(1\leq i\leq n)$). Furthermore,
$I:S_{Is}(Is,T_{h})\rightrightarrows S_{Ic}(Ic,(T_{h},T_{h}+\tau))$ (3.5)
represents the enabling mapping from the source state $S_{Is}(Is,T_{h})$ to the channel state $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$. Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{Is}(Is,T_{h})$ and the complex implementation mechanisms of $S_{Ic}(Ic,(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system. The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
3.2 Noisy Information
The realization of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
Definition 3 (Noisy Information) For information $I=\left\langle o,T_{h},S_{o},c,T_{m},S_{c}\right\rangle$, let $\widetilde{S_{o}}\left(o,T_{h}\right)$ be another state of the ontology $o$ at time $T_{h}$, and $\widetilde{S_{c}}\left(c,T_{m}\right)$ be another state of the carrier $c$ at time $T_{m}$,
@@ 3.2 Noisy Information
where the state $S_{o}^{-}\left(o,T_{h}\right)$ is the loss state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}^{-}\left(o,T_{h}\right)\subseteq S_{o}\left(o,T_{h}\right)$ and not being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$, and the state $S_{o}^{+}\left(o,T_{h}\right)$ is the superposed state of $S_{o}\left(o,T_{h}\right)$, satisfying $S_{o}\left(o,T_{h}\right){\cap S}_{o}^{+}\left(o,T_{h}\right)=\varnothing$ and being enabled to map to $\widetilde{S_{c}}\left(c,T_{m}\right)$. The enabling mapping is then defined as:
$\widetilde{I}:\widetilde{S_{o}}\left(o,T_{h}\right)\rightrightarrows\widetilde{S_{c}}\left(c,T_{m}\right)$ (3.7)
which is the noisy information of $I$, while $I$ is its own noiseless information. Thus, the noisy information $\widetilde{I}$ is also information. It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same. Fig.1(b) illustrates the structural characteristics of noisy information.
In fact, Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario (Shannon, 1948). In this case, he assumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
$E=f(S,N)$ (3.8)
where $f$ is the distortion function of the channel to the transmitted signal. According to OIT's view, the signal $S$ here is the ontological state of information $I$, $S_{Is}\left(Is,T_{h}\right)$. The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{Is}^{-}\left(Is,T_{h}\right)$ from the source state $S_{Is}\left(Is,T_{h}\right)$ and adding another state $S_{Is}^{+}\left(Is,T_{h}\right)$. The output $E$ is then the channel's output state $\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output. Thus,
$\widetilde{I}:\widetilde{S_{Is}}\left(Is,T_{h}\right)\rightrightarrows\widetilde{S_{Ic}}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.9)
is the noisy information of
$I:S_{Is}\left(Is,T_{h}\right)\rightrightarrows S_{Ic}\left(Ic,(T_{h},T_{h}+\tau)\right)$ (3.10)